OpenStudy (anonymous):

What is the product of the product of 3 times the square root of 6 x and 2 times the square root of 15 x? Simplify if possible. 54xsquare root of 10 8x2square root of 10 18xsquare root of 10 18square root of the product of 10 and x squared

OpenStudy (anonymous):

can someone please help im doing my final exam and ive been working on this exam for the last two hours

OpenStudy (anonymous):

Try splitting up into parentheses so it's more clear. I'm not 100% sure what your problem is supposed to look like; is it like this? ... the product of [the product of 3 * times sqrt(6x) AND 2 * sqrt(15x)] \[3*\sqrt{6x} * 2*\sqrt{15x}\] Is that what it is? It might help if you used the 'Draw' tool at the bottom of this comment box if what I typed out is not what you meant.

OpenStudy (anonymous):

idk cuz its multiple choice im just so tires ive been doing this exam for 2 hours and there were like 80 questions

OpenStudy (zzr0ck3r):

multiplication is associative so just multiply all those numbers together

OpenStudy (anonymous):

i just need someone to choose from the multiple choice...thanks

OpenStudy (zzr0ck3r):

we are not here to take your test

OpenStudy (anonymous):

We don't just give out answers on here -- the people on this website want to help students learn by guiding them through the steps. So if you really want the answer then you will just have to work through it and we can help with that.

OpenStudy (zzr0ck3r):

hint sqrt(6x)*sqrt(15x) = sqrt(90x) = sqrt(9*10x) = sqrt(9)*sqrt(10x)

OpenStudy (zzr0ck3r):

so you had 3*2(sqrt(6x)*sqrt(15x) now you have 3*2(sqrt(9)*sqrt(10)) = 3*2*sqrt(9)*sqrt(10)

OpenStudy (zzr0ck3r):

another hint sqrt(9) = 3